OpenStudy (caroline1218):

Which equation has a graph perpendicular to the graph of 7x = 14y - 8 ?

OpenStudy (caroline1218):

y=-2x-7 y=-1/2x+4 y=1/2x+1 y=2x+9

OpenStudy (caroline1218):

@imqwerty @nincompoop @pooja195 @Thadds2003

OpenStudy (thadds2003):

Okay, to the first item, you need to know the way to build a line equation with the slope and a point.

OpenStudy (caroline1218):

@Thadds2003 So then y = 1/2x + 4/7

OpenStudy (caroline1218):

@Thadds2003

OpenStudy (thadds2003):

Yes, im pretty sure.

OpenStudy (thadds2003):

Sorry I cant help any longer, I got to go.

OpenStudy (caroline1218):

Ok

OpenStudy (caroline1218):

@Decarr432 Do you know which equation is perpendicular to y=1/2x+4/7 ?

OpenStudy (decarr432):

I normally use a graphing calculator sorry I won't be of any help to learn

OpenStudy (lynfran):

we first find he slope of the given line 7x=14y-8

OpenStudy (caroline1218):

I believe the answer is A. Is this correct?

OpenStudy (lynfran):

14y-7x-8=0 solve for y we get 14y=7x+8 y=(7/14)x+8/14 which reduce to. y=1/2x+8/14 so the equation of the line is in the form y=mx+b where m=slope our slope is 1/2 to find its perpendicular slope, the product of the two slopes must equals -i m*m1=-1 (1/2)*m1=-1 m1=??

OpenStudy (lynfran):

yes ur answer choice is correct

OpenStudy (caroline1218):

Ok, thanks :)